OpenStudy (anonymous):

For what values of a and b would (a+b)^2 = a^2 + b^2 ? Prove your answer

OpenStudy (shubhamsrg):

when ab =0 ,,the above will hold true..

OpenStudy (shubhamsrg):

i.e. either a=0 or b=0 or both =0

OpenStudy (shubhamsrg):

just expand (a+b)^2 and you should get your ans.

Parth (parthkohli):

f(a or b) = a^2 + b^2 Only x = 0 would prove this. f(0) = 0^2 + 0^2 f(0) = 0

Parth (parthkohli):

I mean a or b = 0, not x = 0

OpenStudy (anonymous):

Thanks @shubhamsrg and @ParthKohli :))))